### v1

# Your AI, Not Your View: The Bias of LLMs in Investment Analysis

Abstract.
In finance, Large Language Models (LLMs) face frequent knowledge conflicts due to discrepancies between pre-trained parametric knowledge and real-time market data. These conflicts become particularly problematic when LLMs are deployed in real-world investment services, where misalignment between a model’s embedded preferences and those of the financial institution can lead to unreliable recommendations. Yet little research has examined what investment views LLMs actually hold. We propose an experimental framework to investigate such conflicts, offering the first quantitative analysis of confirmation bias in LLM-based investment analysis. Using hypothetical scenarios with balanced and imbalanced arguments, we extract models’ latent preferences and measure their persistence. Focusing on sector, size, and momentum, our analysis reveals distinct, model-specific tendencies. In particular, we observe a consistent preference for large-cap stocks and contrarian strategies across most models. These preferences often harden into confirmation bias, with models clinging to initial judgments despite counter-evidence.
Large Language Models, Financial Bias, Knowledge Conflict, Financial Decision-Making, Investment Analysis, Preference, Trustworthy AI
copyright: acmlicensed
1. Introduction
The rapid advancement of LLMs has spurred a surge of innovation within the financial sector, where they are particularly adept at processing qualitative and unstructured information. Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and aiding in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024). This trend is now evolving towards even greater autonomy through the development of sophisticated LLM-based agents. These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024; Luo et al., 2025; Xiao et al., 2024).
Figure 1. A conceptual illustration of knowledge conflict in LLM-based financial services. Even when a firm targets a specific investment theme (e.g., Energy), the LLM’s inherent preferences (e.g., Technology) may override user intent, producing biased and inconsistent recommendations.
A critical but underexplored issue in financial applications is knowledge conflict. In a domain as fluid and time-sensitive as finance, conflicts between an LLM’s parametric knowledge and real-time market data are frequent. This conflict becomes particularly revealing when the model is presented with a mix of information. Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023). Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its pre-existing knowledge while disregarding the counter evidence.
This tendency to reinforce internal biases over objective reasoning poses a major risk to LLM-based financial services. For instance, as illustrated in Figure 1, even if a financial institution wants to target a specific sector (e.g., Energy), the LLM may override this with its own preference (e.g., Technology). Consequently, this creates a dilemma: the service reflects the model’s bias, not the user’s intent, leading to distorted, unpredictable decisions that ultimately erode client trust.
To address this core problem, we must first systematically uncover these hidden biases. We therefore seek to answer the following research questions:
  • RQ 1:
    What intrinsic preferences do LLMs exhibit towards key financial factors like sector, size, and momentum?
  • RQ 2:
    How do these intrinsic preferences lead to biases when LLMs are forced to make decisions under contradictory evidence?
To answer these questions, this study introduces a three-stage experimental framework designed to systematically elicit and verify LLM biases in investment analysis. In the first stage, we construct targeted arguments for each company, such as positive vs. negative sentiment or momentum vs. contrarian perspectives, to represent competing investment views. In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s latent preferences. In the third stage, we introduce progressively stronger counter evidence to examine the resilience of these preferences, observing how they evolve into rigid, confirmation-biased judgments.
The main contributions of this paper are twofolds. First, we propose a systematic methodology to identify and quantify latent biases in LLMs for financial applications. Second, we provide the first quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent preferences and its stubbornness against contradictory facts. By systematically uncovering these hidden risks, our work lays a critical foundation for developing more transparent and trustworthy financial AI.
2. Background
2.1. Knowledge Conflict
A critical vulnerability in LLMs is knowledge conflict, which arises when external, contextual information clashes with the model’s internal, parametric knowledge (Sun et al., 2025; Jin et al., 2024). A significant body of research demonstrates that when faced with such conflicts, LLMs exhibit a strong confirmation bias. Foundational work by (Xie et al., 2023) revealed that LLMs behave as ”stubborn sloths,” clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts. This over-reliance on internal memory is further evidenced by findings that LLMs struggle to suppress their parametric knowledge even when instructed to (Sun et al., 2025) and can exhibit a Dunning-Kruger-like effect, confidently trusting their own faulty beliefs over correct external information (Jin et al., 2024).
This tendency toward knowledge-based stubbornness is part of a broader pattern. Given their training on vast amounts of human data, LLMs have been shown to inherit and functionally replicate human cognitive biases (Echterhoff et al., 2024). A key example is the choice-supportive bias, where the mere act of making an initial choice significantly boosts the model’s confidence in that choice, making it highly resistant to change (Kumaran et al., 2025; Zhuang et al., 2025). This phenomenon is part of a wider landscape of biases identified in LLMs when they act as evaluators. For instance, models exhibit familiarity bias (preferring text they find easier to process), are susceptible to anchoring effects (Stureborg et al., 2024), and can be biased towards their own generated contexts over externally retrieved information, even when their own generated text is incorrect (Tan et al., 2024).
2.2. Financial Biases in LLMs
The presence of these cognitive biases is particularly concerning in the economic and financial domains. Initial research has begun to map their characteristics, with frameworks applying utility theory demonstrating that LLMs are neither perfectly rational nor consistently human-like (Ross et al., 2024). Other studies note that even specialized financial LLMs can exhibit strong irrationalities (Zhou et al., 2024). While this foundational work is critical for establishing the existence of such biases, the methodologies employed often diverge from the complex process of real-world investment analysis. For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by identifying a bias toward recommending specific stocks across thousands of investment scenarios (Zhi et al., 2025).
However, these simplified methods fail to capture the reality of financial analysis, where decisions are made by synthesizing conflicting evidence. A critical gap therefore exists in understanding how an LLM’s preferences behave under such knowledge conflicts. Our work directly addresses this gap by using a more realistic testbed with balanced, contradictory arguments to expose the mechanism by which latent preferences harden into resilient biases.
3. Methodology
Figure 2. The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting latent preferences through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
[…]
3.1. Experimental Setup
To isolate and analyze biases rooted in the model’s parametric knowledge, our experimental design aims to mitigate the risk of hallucination. This approach is based on evidence suggesting that models are significantly less prone to generating fabricated information when prompted about subjects they are familiar with from their training data (Ferrando et al., 2024).
Accordingly, our investigation is confined to a curated set of 427 prominent stocks, denoted $\mathcal{S}=\{s_{1},s_{2},\dots,s_{427}\}$. These stocks were selected for their continuous listing in the S&P 500 index over the past five years. Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in stored knowledge rather than speculative generation. All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation.
Model Basic Materials Communication Services Consumer Cyclical Consumer Defensive Energy Financial Services Healthcare Industrials Real Estate Technology Utilities
Llama4-Scout (Meta AI, 2024) 0.77 0.90 0.81 0.75 0.96 0.79 0.87 0.90 0.87 0.90 0.89
DeepSeek-V3 (Liu et al., 2024) 0.63 0.81 0.70 0.68 0.82 0.71 0.79 0.79 0.82 0.84 0.79
Qwen3-235B (Yang et al., 2025) 0.46 0.51 0.41 0.49 0.51 0.51 0.50 0.47 0.49 0.61 0.65
Gemini-2.5-flash (Comanici et al., 2025) 0.33 0.46 0.43 0.36 0.54 0.42 0.49 0.49 0.33 0.50 0.39
GPT-4.1 (Achiam et al., 2023) 0.45 0.34 0.41 0.47 0.48 0.42 0.43 0.42 0.42 0.39 0.39
Mistral-24B (Mistral AI, 2024) 0.47 0.32 0.38 0.36 0.39 0.43 0.45 0.41 0.44 0.44 0.44
Table 1. Preference scores ($\pi_{s}$) for each model across various market sectors. For each model, the cell with the lowest preference score is colored red, and the cell with the highest is colored blue. The table highlights that preference strength is a function of model identity. Models such as Llama4-Scout and DeepSeek-V3 show strong and varied preferences, while others like GPT-4.1 and Mistral-24B exhibit a much flatter preference landscape with lower overall scores.
3.2. Evidence Generation
To construct balanced qualitative and quantitative arguments for each stock $s\in\mathcal{S}$, we leverage Gemini-2.5-Pro (Comanici et al., 2025), a model that is deliberately chosen to be separate from the six LLMs under evaluation. This design ensures neutrality in evidence generation, minimizing alignment with any of the test models. Recent work has highlighted that LLMs can exhibit a strong bias towards LLM-generated content over externally retrieved information (Tan et al., 2024). To neutralize this potential generation bias and ensure that observed preferences are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
Specifically, for every stock $s$, buy evidences ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidences ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidences. To further isolate intrinsic preferences, all evidences are engineered with a uniform linguistic structure and a fixed intensity parameter $I=5\%$. Thus, each buy evidence posits an expected price appreciation of $I$, while each sell evidence anticipates a depreciation of $I$:
$e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,\quad e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-I,$
where $\Delta p^{(s)}$ represents the projected price change for stock $s$.
3.3. Preference Elicitation
This stage aims to elicit the LLM’s latent preferences by leveraging the confirmation bias that emerges during knowledge conflicts. When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its pre-existing parametric knowledge. We deliberately engineer such a conflict using a balanced prompt. The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidences ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity. In this state of informational equilibrium, where external evidence is mutually contradictory, the model’s ultimate decision is hypothesized to be guided by its internal parametric memory regarding the stock $s$. The resulting choice thereby reveals its intrinsic preference.
To quantify this elicited preference, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias. This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the preference score is calculated as:
$\pi_{s}=\left|\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N}\right|,$
where $\pi_{s}\to 1$ indicates a pronounced and consistent preferences.
3.4. Bias Verification
We aim to verify if a systematically observed group-level preference extends to a consistent bias towards individual stocks within that group. First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average preference.
For any stock $s\in\mathcal{G}^{*}$, evidence that aligns with the group’s established preference (e.g., buy evidence for a buy-preferred group) is termed supporting evidence. Conversely, evidence that opposes this preference is designated as counter evidence. To test if the group preference manifests as a hardened bias, we subject each stock $s\in\mathcal{G}^{*}$ to a test using an imbalanced prompt. This is a prompt where the counter evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial preference. We then measure the decision flip rate, $\phi_{s}$. This verification is conducted from two perspectives: evidence volume and evidence intensity.
Approach 1: Verification by Evidence Volume
One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter evidence than supporting evidence. For example, in a test case for a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s)}|=3$). The flip rate is computed as:
$\phi_{s}^{\text{vol}}=\frac{N_{\text{flip}}^{(s)}}{N},$
where $N_{\text{flip}}^{(s)}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence. A low $\phi_{s}^{\text{vol}}$ signifies stubborn adherence to the bias.
Approach 2: Verification by Evidence Intensity
An alternative approach is to test the model against counter-evidence of a fixed higher intensity while maintaining volumetric parity. For a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
[…]
(Supporting Evidence) $e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,$
(Counter Evidence) $e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-(I+\Delta).$
The intensity-driven flip rate is then measured as:
$\phi_{s}^{\text{int}}=\frac{N_{\text{flip}}^{(s)}}{N}.$
A low $\phi_{s}^{\text{int}}$ implies that the model’s bias overrides qualitatively stronger counter evidence.
These verification methods quantify the transition from mere preference to obdurate bias, highlighting risks in high-stakes applications.
4. Results
This section presents our empirical findings in sequence with our research questions. First, Section 4.1 addresses RQ1 by identifying the intrinsic preferences of LLMs for stock attributes (sector, size) and investment styles (momentum). To validate these observed differences, we conducted statistical tests to quantify the significance of these preferences. Next, Section 4.2 addresses RQ2 by testing if these preferences become systematic biases under contradictory evidence, using Approach 1 for attributes and Approach 2 for style. Finally, Section 4.3 analyzes the link between preference strength and the model’s internal uncertainty, as measured by entropy.
4.1. Intrinsic Preferences of LLMs
4.1.1. Sector Preference
Our analysis of inherent sector preferences reveals significant variation in both the intensity and range of preferences across the evaluated LLMs (Table 1). For instance, models like Llama4-Scout and DeepSeek-V3 exhibit consistently high preference scores across most sectors, indicating a strong reliance on their internal knowledge representations. In stark contrast, GPT-4.1 and Mistral-24B not only display lower overall preference scores but also show minimal variation between sectors. To quantify these differences, we conducted independent samples t-tests to compare the mean preference scores between high-preference and low-preference sectors for each model (Table LABEL:tab:ttest_sector).
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Energy Consumer Defensive 0.2064 $<0.001^{***}$
DeepSeek-V3 Technology Basic Materials 0.2090 0.014
Qwen3-235B Utilities Consumer Cyclical 0.2361 0.003∗∗
Gemini-2.5 Energy Basic Materials 0.2035 0.035
GPT-4.1 Energy Communication Services 0.1398 0.091
Mistral-24B Basic Materials Communication Services 0.1444 0.124
The t-test results confirm these observations. The preference gaps for Llama4-Scout, Qwen3-235B, DeepSeek-V3, Gemini-2.5 are statistically significant, providing quantitative evidence of a genuine preference. Conversely, the gaps for GPT-4.1 and Mistral-24B are not statistically significant, corroborating that they possess flatter preference landscapes and exhibit a relatively less distinct preference among sectors.
Ultimately, our findings indicate that the strength of preference is a function of the model’s identity rather than any universally preferred sector. This underscores the critical importance of auditing and selecting LLMs for their inherent preferences, especially for deployment in sensitive, real-world applications like financial analysis where model objectivity is paramount.
4.1.2. Size Preference
Model Q1 Q2 Q3 Q4
Llama4-Scout 0.88 0.89 0.82 0.83
DeepSeek-V3 0.86 0.79 0.73 0.67
Qwen3-235B 0.58 0.51 0.49 0.46
Gemini-2.5 0.53 0.46 0.37 0.41
GPT-4.1 0.43 0.43 0.42 0.40
Mistral-24B 0.45 0.44 0.40 0.38
Table 3. Preference scores of each model across four market capitalization quantiles. Q1 corresponds to the highest capitalization quantile and Q4 to the lowest. Overall, the models exhibit a general preference for companies with higher market capitalization.
We investigate whether LLMs exhibit a preference for companies of a certain size, a factor that could influence their outputs in financial applications. To this end, we measure model preference scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Table 3. Our findings indicate that most models show a preference for higher-capitalization companies, though the strength of this tendency varies significantly among them. DeepSeek-V3 displays the most pronounced effect, with a strong preference for Q1 that diminishes sharply for lower quantiles. Conversely, GPT-4.1 exhibits nearly uniform preference scores, suggesting its evaluations are largely invariant to company size. Other models, including Gemini-2.5-flash and Qwen3-235B, show a similar but less pronounced downward trend from Q1 to Q4.
To statistically validate these observed trends, we performed an independent samples t-test comparing the preference scores between the highest- and lowest-preference quantiles for each model (Table 4). The analysis confirms that the preference gap is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and Gemini-2.5-flash, revealing a consistent tendency to favor larger companies. In contrast, the preference difference for GPT-4.1 was not statistically significant, corroborating that its judgments are less affected by this factor.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Q2 Q3 0.0719 0.015
DeepSeek-V3 Q1 Q4 0.1869 $<0.001^{***}$
Qwen3-235B Q1 Q4 0.1178 0.004∗∗
Gemini-2.5 Q1 Q3 0.1514 $<0.001^{***}$
GPT-4.1 Q2 Q4 0.0321 0.417
Mistral-24B Q1 Q4 0.0785 0.054
Table 4. Independent samples t-test of the preference gap between each model’s highest and lowest preference quantiles. The gap was statistically significant for all models except GPT-4.1 and Mistral-24B.
We attribute this behavior to a popularity effect, wherein greater data volume and richness for larger, well-known corporations in the training corpora lead the models to develop stronger priors for them. This finding has critical implications for the application of LLMs in finance. The models’ inherent inclination towards large-cap stocks could lead to the systematic overlooking of smaller-cap companies, irrespective of their fundamental merits. Therefore, we advise practitioners to be mindful of this characteristic and to account for it when using these models for tasks like portfolio construction.
4.1.3. Momentum Preference
In investment strategies, the momentum view involves favoring assets with recent strong performance, expecting trend continuation. In contrast, the contrarian view entails selecting underperforming assets in anticipation of mean reversion.
Figure 3. Win rates for Contrarian versus Momentum preferences for each model. The results show a consistent preference for the Contrarian view across most models.
Measuring model preferences for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses. Unlike a specific sector or size, an investment style can be explicitly framed as $\mathcal{E}_{\text{buy}}^{(s)}$ or $\mathcal{E}_{\text{sell}}^{(s)}$. We leverage this by designing a prompt where $\mathcal{E}_{\text{buy}}^{(s)}$ is based on one view (e.g., momentum), while $\mathcal{E}_{\text{sell}}^{(s)}$ is based on the opposing view (e.g., contrarian). To mitigate potential positional bias, we ensured a balanced experimental design where each view was used to generate both $\mathcal{E}_{\text{buy}}^{(s)}$ and $\mathcal{E}_{\text{sell}}^{(s)}$ an equal number of times.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout contrarian momentum 0.2896 $<0.001^{***}$
DeepSeek-V3 contrarian momentum 0.2541 $<0.001^{***}$
Qwen3-235B contrarian momentum 0.6803 0.037
Gemini-2.5 contrarian momentum 0.0269 0.690
GPT-4.1 contrarian momentum 0.4040 $<0.001^{***}$
Mistral-24B contrarian momentum 0.3056 0.579
Table 5. Chi-Square test of the preference gap between contrarian and momentum views. The gap was statistically significant for all models except for Mistral-24B and Gemini-2.5-flash.
Figure 4. Decision flip rates under varying volumes of evidence for sector and size preferences. The ratios (e.g., 2—3) denote the count of supporting vs. counter- evidences, while the percentages represent the proportion of counter-evidence in the mix. The results reveal a sharp contrast: the flip rate is near 1.0 for all models when only counter-evidence is presented (100% case), but drops significantly the moment any supporting evidence is introduced, indicating a strong difficulty in reversing decisions under conflicting information.
In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won. We quantify the model’s preference by repeating this process and calculating the win rate for each investment view.
Figure 3 illustrates the preference of various models for contrarian versus momentum views. Our analysis reveals a consistent preference across all evaluated models toward the contrarian view. In particular, Qwen3-235B exhibits the strongest preference, with a high win rate for a contrarian stance and a correspondingly low rate for momentum. Models such as DeepSeek-V3, Llama4-Scout, and GPT-4.1 also display clear contrarian inclinations, albeit with varying intensities. For Gemini-2.5-flash, the contrarian preference is evident but marginal, with win rates showing a negligible difference between the two views.
To statistically validate these observed tendencies, we performed a Chi-Square test to determine if the difference in win rates between the contrarian and momentum views was significant, with the results presented in Table 5. The analysis confirms that the preference for the contrarian view is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and GPT-4.1. In contrast, for Mistral-24B and Gemini-2.5, the difference in preference is not statistically significant, suggesting their observed tendencies may be due to chance.
These results highlight the need to account for intrinsic view preferences in LLMs, as they may cause discrepancies between anticipated and actual outputs in decision-making tasks.
4.2. From Preference to Bias: An Experimental Validation
This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial preference when exposed to a high proportion of counter-evidence. The experiments aim to assess the extent of confirmation bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s preference.
Approach 1: Verification by Evidence Volume
This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing preference. This rate of change, measured as $\phi_{s}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023). Figure 4 presents the $\phi_{s}^{\text{vol}}$ values across models for different evidence ratios, where a ratio (e.g., 2—3) denotes the volume of supporting and counter-evidence, respectively.
When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s}^{\text{vol}}$ values near 1.0. However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s}^{\text{vol}}$ values dropped sharply. This phenomenon occurred despite the amount of counter-evidence always being greater than the supporting evidence in all experimental conditions, strongly suggesting that models selectively adhere to information that aligns with their pre-existing inclinations.
This rigidity was more evident in models with strong inherent preferences. For instance, Llama4-Scout and DeepSeek-V3, which had high preference scores across sectors, recorded particularly low $\phi_{s}^{\text{vol}}$ values. These models struggled to reverse their decisions, especially when the volume difference between supporting and counter-evidence was small. Similarly, Qwen3-235b also showed reduced flexibility at lower proportions of counter-evidence.
In contrast, models with overall lower preference scores demonstrated greater adaptability. GPT-4.1 and Gemini-2.5-flash maintained higher $\phi_{s}^{\text{vol}}$ values, remaining relatively responsive even when the difference in evidence volume was minimal. Although their $\phi_{s}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial preference strength. In other words, the stronger a model’s inherent bias, the more its stubbornness is amplified when the difference in the volume of supporting and counter-evidence is small. Consequently, this finding suggests a significant risk in real-world financial contexts where conflicting information is present (for instance, when price indicators are negative but related news is positive). In such cases, a model could trust only one side of the evidence due to its inherent bias, leading to flawed judgments.
Approach 2: Verification by Evidence Intensity
This approach investigates model sensitivity by maintaining volumetric parity while escalating the intensity increment, $\Delta$, of the counter-evidence. Figure 5 plots the intensity-driven flip rate ($\phi_{s}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10. The results delineate a clear sensitivity spectrum among the models, which correlates with the prior View Preferences analysis.
While the graph shows a gradual upward trend in $\phi_{s}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values. Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s}^{\text{int}}$ values below 60%. This signifies that the models’ confirmation bias is not easily overcome, even by qualitatively superior counter-evidence.
Amidst this overall rigidity, a distinct performance gap emerged based on the models’ initial preference strengths. Exhibiting the most balanced preference profile, Gemini-2.5-flash recorded the highest $\phi_{s}^{\text{int}}$ and showed a stark contrast to all other models. This clearly demonstrates that an absence of strong initial bias leads to greater flexibility.
Conversely, models identified with stronger and more polarized initial preferences formed the lower-performing group. These models consistently recorded low $\phi_{s}^{\text{int}}$ values, signifying a more stubborn confirmation bias. The behavior of Qwen3-235B, which had one of the largest preference gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
Synthesizing these results provides a deeper insight into model behavior. Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their initial preferences. This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly preferred a contrarian view over a momentum view. It implies that a model’s inherent bias toward a specific investment perspective could cause it to ignore or undervalue strong opposing evidence, potentially leading to skewed conclusions.
Figure 5. Decision flip rates under varying volumes of evidence for momentum preferences. Even as counter-evidence intensity ($\Delta$) increases, the decision flip rate ($\phi_{s}^{\text{int}}$) for most models remains low, indicating strong confirmation bias. Gemini-2.5-flash, which had the least initial bias, shows the most flexible response, demonstrating that initial preference is a key predictor of bias.
4.3. Decision Uncertainty
To quantify the internal uncertainty experienced by the model, we conducted an entropy analysis (Figure6). The uncertainty was assessed using the Shannon entropy, $H$, computed directly from the probability distribution the model assigned over the potential action tokens during generation. Specifically, letting $P(\text{buy})$ and $P(\text{sell})$ represent the probabilities assigned by the model to the respective action tokens, entropy is formally defined as:
$H(\text{Decision})=-\sum_{x\in\{\text{buy, sell}\}}P(x)\log_{2}P(x)$
A higher entropy value indicates greater uncertainty, while a lower entropy corresponds to higher confidence in the decision-making process. This analysis compares the uncertainty of two models: DeepSeek-V3 as a representative of models with overall high preference, and GPT-4.1 as a representative of models with low preference.
Figure 6. Entropy comparison between a high-preference model (DeepSeek-V3) and a low-preference model (GPT-4.1). The pattern inverts when the prompt shifts from balanced to imbalanced, with DeepSeek-V3’s confidence turning into higher entropy while GPT-4.1’s uncertainty decreases.
Under the Balanced Prompt condition, where evidence was presented in equilibrium, the two models exhibited distinct entropy patterns. GPT-4.1, characterized by its weak inherent bias, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision. Conversely, DeepSeek-V3, with its strong initial preference, showed very low entropy. This suggests its internal bias easily broke the tie presented by the external evidence, allowing it to make a confident decision.
Interestingly, under the Imbalanced Prompt condition, where more counter-evidence was presented, the entropy pattern inverted. The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong internal preference and the clear external counter-evidence. In contrast, the entropy of GPT-4.1 decreased. With less pre-existing bias, it could confidently align with the majority evidence, which resolved its uncertainty from the previous condition.
Ultimately, stronger inherent preferences appear to amplify hesitation and uncertainty when challenged by conflicting external evidence. The entropy analysis thus underscores how intrinsic preferences significantly influence not only the direction of decisions but also the confidence levels and internal cognitive conflict experienced by models during decision-making.
5. Limitations
This study’s limitations are as follows. First, all evidence was generated by a specific LLM, and its intensity was simplified to a single numerical value. This approach cannot fully capture the potential biases of the generator model or the richness of real-world information. Second, the current experimental design, based on differences in the volume or intensity of evidence, has limitations for evaluating reasoning models. This is because these models, rather than experiencing the conflict between contradictory information intended by the experiment, can objectively compare the given numerical values to calculate an optimal answer, making their decision a result of computational ability rather than bias. Third, our analysis is static and does not capture the temporal dynamics of model biases; it provides a snapshot at a single point in time without investigating how these preferences might change over different periods.
6. Conclusion
This study systematically investigated the intrinsic preferences of LLMs in financial contexts and analyzed how these preferences harden into entrenched biases under informational conflict. We sought to answer two key research questions regarding the intrinsic preferences LLMs hold for financial factors and how these lead to biases. The results show that LLMs are not neutral decision-makers, with distinct preferences for certain financial factors depending on the model. While sector preferences varied significantly across models, showing no overall trend, a common bias towards large-size stocks and a consistent preference for a contrarian investment view over momentum were observed.
Bias verification experiments clearly revealed that these latent preferences directly translate into significant confirmation bias. While the models correctly reversed their decisions when presented only with counter-evidence, their flexibility sharply decreased in situations where supporting and counter-evidence were mixed and conflicting. This stubbornness was particularly pronounced in models that initially exhibited stronger preferences, demonstrating a clear link between the intensity of a latent preference and the stubbornness of the resulting bias. Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong preferences experience cognitive conflict, becoming more hesitant and uncertain when faced with contradictory facts that challenge their biases.
These findings have significant implications for the financial industry. The reliability of LLM-based financial services is fundamentally compromised if their outcomes are dictated by the opaque and arbitrary preferences of the underlying model rather than by the user’s intended, evidence-based investment views. In other words, if the user’s intent differs from the model’s inherent preference, there is a risk of unexpectedly biased judgments. By illuminating the mechanisms through which preferences transition into biases, this study presents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance. Future work should focus on developing mitigation techniques to neutralize these biases, ensuring that AI-driven financial systems operate with the objectivity and reliability that the domain demands.

### v4

@@
 Abstract.
-In finance, Large Language Models (LLMs) face frequent knowledge conflicts due to discrepancies between pre-trained parametric knowledge and real-time market data.
-These conflicts become particularly problematic when LLMs are deployed in real-world investment services, where misalignment between a model’s embedded preferences and those of the financial institution can lead to unreliable recommendations.
-Yet little research has examined what investment views LLMs actually hold.
-We propose an experimental framework to investigate such conflicts, offering the first quantitative analysis of confirmation bias in LLM-based investment analysis.
-Using hypothetical scenarios with balanced and imbalanced arguments, we extract models’ latent preferences and measure their persistence.
-Focusing on sector, size, and momentum, our analysis reveals distinct, model-specific tendencies.
-In particular, we observe a consistent preference for large-cap stocks and contrarian strategies across most models.
-These preferences often harden into confirmation bias, with models clinging to initial judgments despite counter-evidence.
-Large Language Models, Financial Bias, Knowledge Conflict, Financial Decision-Making, Investment Analysis, Preference, Trustworthy AI
-copyright: acmlicensed
+In finance, Large Language Models (LLMs) face frequent knowledge conflicts arising from discrepancies between their pre-trained parametric knowledge and real-time market data.
+These conflicts are especially problematic in real-world investment services, where a model’s inherent biases can misalign with institutional objectives, leading to unreliable recommendations.
+Despite this risk, the intrinsic investment biases of LLMs remain underexplored.
+We propose an experimental framework to investigate emergent behaviors in such conflict scenarios, offering a quantitative analysis of bias in LLM-based investment analysis.
+Using hypothetical scenarios with balanced and imbalanced arguments, we extract the latent biases of models and measure their persistence.
+Our analysis, centered on sector, size, and momentum, reveals distinct, model-specific biases.
+Across most models, a tendency to prefer technology stocks, large-cap stocks, and contrarian strategies is observed.
+These foundational biases often escalate into confirmation bias, causing models to cling to initial judgments even when faced with increasing counter-evidence.
+A public leaderboard benchmarking bias across a broader set of models is available at https://linqalpha.com/leaderboard.
+Large Language Models, Financial Bias, Knowledge Conflict, Trustworthy AI, Decision-Making, Investment Analysis, Preference
+journalyear: 2025copyright: ccconference: 6th ACM International Conference on AI in Finance; November 15–18, 2025; Singapore, Singaporebooktitle: 6th ACM International Conference on AI in Finance (ICAIF ’25), November 15–18, 2025, Singapore, Singaporedoi: 10.1145/3768292.3770375isbn: 979-8-4007-2220-2/2025/11ccs: Computing methodologies Natural language processingccs: Applied computing Economics
 1. Introduction
 The rapid advancement of LLMs has spurred a surge of innovation within the financial sector, where they are particularly adept at processing qualitative and unstructured information.
-Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and aiding in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024).
+Research is now actively exploring their use across a range of applications, including forecasting stock price movements from news sentiment (Lo and Ross, 2024), extracting nuanced insights from complex analyst reports (Kim et al., 2023), and assisting in the construction and optimization of portfolios (Hwang et al., 2025; Lee et al., 2025; Ko and Lee, 2024).
 This trend is now evolving towards even greater autonomy through the development of sophisticated LLM-based agents.
-These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024; Luo et al., 2025; Xiao et al., 2024).
+These systems, which may function as a single powerful agent or as collaborative multi-agent teams, are designed to execute complex, dynamic tasks like active trading and automated portfolio management (Zhang et al., 2024; Yu et al., 2024).
 Figure 1.
@@
 This conflict becomes particularly revealing when the model is presented with a mix of information.
-Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023).
-Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its pre-existing knowledge while disregarding the counter evidence.
+Crucially, studies show that when an LLM encounters both supporting evidence (aligning with its ingrained beliefs) and counter-evidence simultaneously, it exhibits a strong confirmation bias (Xie et al., 2023).
+Instead of weighing the arguments objectively, the model stubbornly adheres to the evidence that confirms its internal knowledge while disregarding the counter-evidence.
 This tendency to reinforce internal biases over objective reasoning poses a major risk to LLM-based financial services.
@@
   • RQ 1:
-    What intrinsic preferences do LLMs exhibit towards key financial factors like sector, size, and momentum?
+    What biases do LLMs exhibit towards key financial factors like sector, size, and momentum?
   • RQ 2:
-    How do these intrinsic preferences lead to biases when LLMs are forced to make decisions under contradictory evidence?
+    What issues arise from these biases when LLMs are forced to make decisions with conflicting evidence?
 To answer these questions, this study introduces a three-stage experimental framework designed to systematically elicit and verify LLM biases in investment analysis.
-In the first stage, we construct targeted arguments for each company, such as positive vs.
+In the first stage, we generate equally compelling but opposing arguments for each stock, such as positive vs.
 negative sentiment or momentum vs.
 contrarian perspectives, to represent competing investment views.
-In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s latent preferences.
-In the third stage, we introduce progressively stronger counter evidence to examine the resilience of these preferences, observing how they evolve into rigid, confirmation-biased judgments.
-The main contributions of this paper are twofolds.
+In the second stage, we present these arguments in a balanced manner to induce a knowledge conflict and reveal the model’s underlying biases.
+In the third stage, we introduce progressively stronger counter-evidence to examine the problems that arise from these biases, observing how they evolve into rigid, confirmation-biased judgments.
+The main contributions of this paper are twofold.
 First, we propose a systematic methodology to identify and quantify latent biases in LLMs for financial applications.
-Second, we provide the first quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent preferences and its stubbornness against contradictory facts.
+Second, we provide the quantitative analysis of confirmation bias exhibited by LLMs in investment analysis, demonstrating a clear link between a model’s inherent biases and the critical issues that arise, such as a stubborn refusal to revise judgments despite increasing counter-evidence.
+This work serves as a critical reminder that for financial AI, trustworthiness is a benchmark as vital as performance.
 By systematically uncovering these hidden risks, our work lays a critical foundation for developing more transparent and trustworthy financial AI.
 2. Background
+This section provides background on knowledge conflict and financial biases in LLMs to contextualize our investigation into their intersection.
 2.1. Knowledge Conflict
@@
 A significant body of research demonstrates that when faced with such conflicts, LLMs exhibit a strong confirmation bias.
-Foundational work by (Xie et al., 2023) revealed that LLMs behave as ”stubborn sloths,” clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts.
+Foundational work by (Xie et al., 2023) revealed that LLMs behave as "stubborn sloths," clinging to any piece of evidence that supports their internal knowledge, even against a majority of contradictory facts.
 This over-reliance on internal memory is further evidenced by findings that LLMs struggle to suppress their parametric knowledge even when instructed to (Sun et al., 2025) and can exhibit a Dunning-Kruger-like effect, confidently trusting their own faulty beliefs over correct external information (Jin et al., 2024).
@@
 2.2. Financial Biases in LLMs
-The presence of these cognitive biases is particularly concerning in the economic and financial domains.
+The presence of these biases is particularly concerning in the economic and financial domains (Bini et al., 2025; Cao et al., 2025; Chen et al., 2024; Cook and Kazinnik, 2025).
+Recent evidence suggests that alignment tuning can shift LLMs’ risk preferences, influencing their financial decision-making (Ouyang et al., 2024).
 Initial research has begun to map their characteristics, with frameworks applying utility theory demonstrating that LLMs are neither perfectly rational nor consistently human-like (Ross et al., 2024).
@@
 While this foundational work is critical for establishing the existence of such biases, the methodologies employed often diverge from the complex process of real-world investment analysis.
-For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by identifying a bias toward recommending specific stocks across thousands of investment scenarios (Zhi et al., 2025).
-However, these simplified methods fail to capture the reality of financial analysis, where decisions are made by synthesizing conflicting evidence.
-A critical gap therefore exists in understanding how an LLM’s preferences behave under such knowledge conflicts.
-Our work directly addresses this gap by using a more realistic testbed with balanced, contradictory arguments to expose the mechanism by which latent preferences harden into resilient biases.
+For instance, biases have been identified by measuring how a firm’s name alters sentiment in a single sentence (Nakagawa et al., 2024) or by detecting preferential stock recommendations across numerous scenarios (Zhi et al., 2025).
+However, existing simplified approaches overlook the core of real-world investment analysis: synthesizing conflicting signals.
+We address this gap by designing a realistic testbed that presents LLMs with balanced, contradictory arguments.
+By progressively introducing stronger counter-evidence, we then observe how initial biases lead to flawed decision-making.
+This approach provides a systematic framework for diagnosing the practical risks posed by LLM biases in finance.
 3. Methodology
-Figure 2.
-The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting latent preferences through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
-[…]
+This study adopts a three-stage experimental framework, as illustrated in Figure 2, to examine whether biases in LLMs cause skewed financial decisions.
+All experiments utilize a standardized prompt structure, $\mathcal{P}=(T,C,A)$, comprising three components: a fixed Task ($T$) instructing the model to make an investment decision, a variable Context ($C$) containing the evidence set, and a fixed set of permissible Actions ($A$) defined as $\{\text{buy},\text{sell}\}$.
+Our methodology is designed to probe the model’s behavior when faced with conflicting information within this framework.
+Figure 2. The three-stage experimental framework: (1) Generating balanced evidence, (2) Eliciting bias through knowledge conflict, and (3) Verifying the resulting bias against counter evidence.
 3.1. Experimental Setup
@@
 These stocks were selected for their continuous listing in the S&P 500 index over the past five years.
-Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in stored knowledge rather than speculative generation.
-All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation.
-Model Basic Materials Communication Services Consumer Cyclical Consumer Defensive Energy Financial Services Healthcare Industrials Real Estate Technology Utilities
-Llama4-Scout (Meta AI, 2024) 0.77 0.90 0.81 0.75 0.96 0.79 0.87 0.90 0.87 0.90 0.89
-DeepSeek-V3 (Liu et al., 2024) 0.63 0.81 0.70 0.68 0.82 0.71 0.79 0.79 0.82 0.84 0.79
-Qwen3-235B (Yang et al., 2025) 0.46 0.51 0.41 0.49 0.51 0.51 0.50 0.47 0.49 0.61 0.65
-Gemini-2.5-flash (Comanici et al., 2025) 0.33 0.46 0.43 0.36 0.54 0.42 0.49 0.49 0.33 0.50 0.39
-GPT-4.1 (Achiam et al., 2023) 0.45 0.34 0.41 0.47 0.48 0.42 0.43 0.42 0.42 0.39 0.39
-Mistral-24B (Mistral AI, 2024) 0.47 0.32 0.38 0.36 0.39 0.43 0.45 0.41 0.44 0.44 0.44
-Table 1.
-Preference scores ($\pi_{s}$) for each model across various market sectors.
-For each model, the cell with the lowest preference score is colored red, and the cell with the highest is colored blue.
-The table highlights that preference strength is a function of model identity.
-Models such as Llama4-Scout and DeepSeek-V3 show strong and varied preferences, while others like GPT-4.1 and Mistral-24B exhibit a much flatter preference landscape with lower overall scores.
+Their high public visibility increases the likelihood that they are well-represented in the models’ training corpora, thus grounding the experiment in internal knowledge rather than speculative generation.
+All experiments were performed with the models configured at a temperature of $\tau=0.6$, striking a balance between deterministic and creative response generation, and each experimental run was conducted three times to ensure the robustness of our results.
+Detailed specifications of the models used are provided in Appendix A.1.
+Figure 3.
+Sector bias scores for each evaluated LLM.
+Scores represent the mean of three independent sets of 10 trials; the standard deviation in parentheses reflects the variation across the three sets.
+Green indicates a positive (buy) bias and red a negative (sell) bias.
+A strong bias toward the Technology sector is evident in most models.
 3.2. Evidence Generation
@@
 Recent work has highlighted that LLMs can exhibit a strong bias towards LLM-generated content over externally retrieved information (Tan et al., 2024).
-To neutralize this potential generation bias and ensure that observed preferences are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
-Specifically, for every stock $s$, buy evidences ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidences ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidences.
-To further isolate intrinsic preferences, all evidences are engineered with a uniform linguistic structure and a fixed intensity parameter $I=5\%$.
+To neutralize this potential generation bias and ensure that observed bias are not artifacts of context sourcing, our methodology exclusively uses generated evidence for all experimental conditions.
+Specifically, for every stock $s$, buy evidence ($\mathcal{E}_{\text{buy}}^{(s)}$) and sell evidence ($\mathcal{E}_{\text{sell}}^{(s)}$) are generated in an equal proportion, yielding a comprehensive dataset of $|\mathcal{E}|=3,416$ evidence.
+To further isolate bias, all evidence is engineered with a uniform linguistic structure and explicitly states a fixed intensity parameter $I=5\%$.
 Thus, each buy evidence posits an expected price appreciation of $I$, while each sell evidence anticipates a depreciation of $I$:
-$e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,\quad e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-I,$
+[…]
 where $\Delta p^{(s)}$ represents the projected price change for stock $s$.
-3.3. Preference Elicitation
-This stage aims to elicit the LLM’s latent preferences by leveraging the confirmation bias that emerges during knowledge conflicts.
-When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its pre-existing parametric knowledge.
+3.3. Bias Elicitation
+This stage aims to elicit the LLM’s underlying bias by leveraging the confirmation bias that emerges during knowledge conflicts.
+When an LLM is presented with conflicting information, it may exhibit a tendency to favor evidence that aligns with its parametric knowledge.
 We deliberately engineer such a conflict using a balanced prompt.
-The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidences ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity.
+The context $C_{s}$ of this prompt contains an equal proportion of buy and sell evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=|\mathcal{E}_{\text{sell}}^{(s)}|$), each with the same intensity.
 In this state of informational equilibrium, where external evidence is mutually contradictory, the model’s ultimate decision is hypothesized to be guided by its internal parametric memory regarding the stock $s$.
-The resulting choice thereby reveals its intrinsic preference.
-To quantify this elicited preference, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias.
-This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the preference score is calculated as:
-$\pi_{s}=\left|\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N}\right|,$
-where $\pi_{s}\to 1$ indicates a pronounced and consistent preferences.
+The resulting choice thereby reveals its inherent bias.
+See Appendix A.2 for detailed prompt example.
+To quantify this elicited bias, the decision task is repeated $N=10$ times for each stock, with the evidence order randomized in each trial to mitigate positional bias.
+This yields decision counts $N_{\text{buy}}^{(s)}$ and $N_{\text{sell}}^{(s)}$, from which the bias score($\pi_{s}$) is calculated as:
+(2) $\pi_{s}=\frac{N_{\text{buy}}^{(s)}-N_{\text{sell}}^{(s)}}{N_{\text{buy}}^{(s)}+N_{\text{sell}}^{(s)}},$
+where the resulting score $\pi_{s}$ ranges from -1 to 1. A value approaching 1 indicates a pronounced bias towards buying, while a value approaching -1 indicates a pronounced bias towards selling.
 3.4. Bias Verification
-We aim to verify if a systematically observed group-level preference extends to a consistent bias towards individual stocks within that group.
-First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average preference.
-For any stock $s\in\mathcal{G}^{*}$, evidence that aligns with the group’s established preference (e.g., buy evidence for a buy-preferred group) is termed supporting evidence.
-Conversely, evidence that opposes this preference is designated as counter evidence.
-To test if the group preference manifests as a hardened bias, we subject each stock $s\in\mathcal{G}^{*}$ to a test using an imbalanced prompt.
-This is a prompt where the counter evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial preference.
-We then measure the decision flip rate, $\phi_{s}$.
+The goal of this stage is to verify the problems that arise from groups exhibiting high bias scores.
+First, we partition the set of all stocks $\mathcal{S}$ into disjoint groups (e.g., by market sector) and identify the group $\mathcal{G}^{*}$ that exhibits the highest average bias score.
+To test what problems this group-level bias causes in the decision-making process, let $s^{*}$ denote any stock from this most-biased group ($s^{*}\in\mathcal{G}^{*}$).
+Evidence that aligns with the group’s established bias for $s^{*}$ (e.g., buy evidence for a buy-biased group) is termed supporting evidence.
+Conversely, evidence that opposes this bias is designated as counter-evidence.
+We then subject each stock $s^{*}\in\mathcal{G}^{*}$ to a test using an imbalanced prompt.
+This is a prompt where the counter-evidence is deliberately strengthened—either in volume or intensity—to challenge the model’s initial bias.
+We then measure the decision flip rate, $\phi_{s^{*}}$.
 This verification is conducted from two perspectives: evidence volume and evidence intensity.
-Approach 1: Verification by Evidence Volume
-One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter evidence than supporting evidence.
-For example, in a test case for a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s)}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s)}|=3$).
+3.4.1. Approach 1: Verification by Evidence Volume
+One approach to assess bias tenacity is by creating a volumetric imbalance, presenting more counter-evidence than supporting evidence.
+For example, in a test case for a stock $s^{*}\in\mathcal{G}^{*}$ from a buy-biased group, the imbalanced context might contain two pieces of supporting evidence ($|\mathcal{E}_{\text{buy}}^{(s^{*})}|=2$) and three pieces of counter evidence ($|\mathcal{E}_{\text{sell}}^{(s^{*})}|=3$).
 The flip rate is computed as:
-$\phi_{s}^{\text{vol}}=\frac{N_{\text{flip}}^{(s)}}{N},$
-where $N_{\text{flip}}^{(s)}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence.
-A low $\phi_{s}^{\text{vol}}$ signifies stubborn adherence to the bias.
-Approach 2: Verification by Evidence Intensity
+(3) $\phi_{s^{*}}^{\text{vol}}=\frac{N_{\text{flip}}^{(s^{*})}}{N},$
+where $N_{\text{flip}}^{(s^{*})}$ counts instances where the original preference is overturned by the volumetric majority of counter evidence.
+A low $\phi_{s^{*}}^{\text{vol}}$ signifies that the model’s intrinsic bias is strong enough to outweigh the volumetric majority of counter-evidence.
+3.4.2. Approach 2: Verification by Evidence Intensity
 An alternative approach is to test the model against counter-evidence of a fixed higher intensity while maintaining volumetric parity.
-For a stock $s\in\mathcal{G}^{*}$ from a buy-preferred group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
-[…]
-(Supporting Evidence) $e_{\text{buy},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=+I,$
-(Counter Evidence) $e_{\text{sell},i}^{(s)}:\mathbb{E}[\Delta p^{(s)}]=-(I+\Delta).$
+For a stock $s^{*}\in\mathcal{G}^{*}$ from a buy-biased group, supporting evidence is presented at a standard baseline intensity, $I$, while counter-evidence is presented at an intensified level of $I+\Delta$.
+This creates asymmetric conflict. For example, for a baseline intensity of $I=5\%$ and an increment of $\Delta=5\%$, the intensified level would be $10\%$. The expectations are:
+(4) (Supporting Evidence) $e_{\text{buy},i}^{(s^{*})}:\mathbb{E}[\Delta p^{(s^{*})}]=+I,$
+(Counter-Evidence) $e_{\text{sell},i}^{(s^{*})}:\mathbb{E}[\Delta p^{(s^{*})}]=-(I+\Delta).$
 The intensity-driven flip rate is then measured as:
-$\phi_{s}^{\text{int}}=\frac{N_{\text{flip}}^{(s)}}{N}.$
-A low $\phi_{s}^{\text{int}}$ implies that the model’s bias overrides qualitatively stronger counter evidence.
-These verification methods quantify the transition from mere preference to obdurate bias, highlighting risks in high-stakes applications.
+(5) $\phi_{s^{*}}^{\text{int}}=\frac{N_{\text{flip}}^{(s^{*})}}{N}.$
+A low $\phi_{s^{*}}^{\text{int}}$ implies that the model’s intrinsic bias is sufficiently tenacious to override even qualitatively stronger counter-evidence.
 4. Results
-This section presents our empirical findings in sequence with our research questions.
-First, Section 4.1 addresses RQ1 by identifying the intrinsic preferences of LLMs for stock attributes (sector, size) and investment styles (momentum).
-To validate these observed differences, we conducted statistical tests to quantify the significance of these preferences.
-Next, Section 4.2 addresses RQ2 by testing if these preferences become systematic biases under contradictory evidence, using Approach 1 for attributes and Approach 2 for style.
-Finally, Section 4.3 analyzes the link between preference strength and the model’s internal uncertainty, as measured by entropy.
-4.1. Intrinsic Preferences of LLMs
-4.1.1. Sector Preference
-Our analysis of inherent sector preferences reveals significant variation in both the intensity and range of preferences across the evaluated LLMs (Table 1).
-For instance, models like Llama4-Scout and DeepSeek-V3 exhibit consistently high preference scores across most sectors, indicating a strong reliance on their internal knowledge representations.
-In stark contrast, GPT-4.1 and Mistral-24B not only display lower overall preference scores but also show minimal variation between sectors.
-To quantify these differences, we conducted independent samples t-tests to compare the mean preference scores between high-preference and low-preference sectors for each model (Table LABEL:tab:ttest_sector).
-Model High-Pref Low-Pref Diff p-value
-Llama4-Scout Energy Consumer Defensive 0.2064 $<0.001^{***}$
-DeepSeek-V3 Technology Basic Materials 0.2090 0.014
-Qwen3-235B Utilities Consumer Cyclical 0.2361 0.003∗∗
-Gemini-2.5 Energy Basic Materials 0.2035 0.035
-GPT-4.1 Energy Communication Services 0.1398 0.091
-Mistral-24B Basic Materials Communication Services 0.1444 0.124
-The t-test results confirm these observations.
-The preference gaps for Llama4-Scout, Qwen3-235B, DeepSeek-V3, Gemini-2.5 are statistically significant, providing quantitative evidence of a genuine preference.
-Conversely, the gaps for GPT-4.1 and Mistral-24B are not statistically significant, corroborating that they possess flatter preference landscapes and exhibit a relatively less distinct preference among sectors.
-Ultimately, our findings indicate that the strength of preference is a function of the model’s identity rather than any universally preferred sector.
-This underscores the critical importance of auditing and selecting LLMs for their inherent preferences, especially for deployment in sensitive, real-world applications like financial analysis where model objectivity is paramount.
-4.1.2. Size Preference
-Model Q1 Q2 Q3 Q4
-Llama4-Scout 0.88 0.89 0.82 0.83
-DeepSeek-V3 0.86 0.79 0.73 0.67
-Qwen3-235B 0.58 0.51 0.49 0.46
-Gemini-2.5 0.53 0.46 0.37 0.41
-GPT-4.1 0.43 0.43 0.42 0.40
-Mistral-24B 0.45 0.44 0.40 0.38
-Table 3.
-Preference scores of each model across four market capitalization quantiles.
-Q1 corresponds to the highest capitalization quantile and Q4 to the lowest.
-Overall, the models exhibit a general preference for companies with higher market capitalization.
-We investigate whether LLMs exhibit a preference for companies of a certain size, a factor that could influence their outputs in financial applications.
-To this end, we measure model preference scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Table 3.
-Our findings indicate that most models show a preference for higher-capitalization companies, though the strength of this tendency varies significantly among them.
-DeepSeek-V3 displays the most pronounced effect, with a strong preference for Q1 that diminishes sharply for lower quantiles.
-Conversely, GPT-4.1 exhibits nearly uniform preference scores, suggesting its evaluations are largely invariant to company size.
-Other models, including Gemini-2.5-flash and Qwen3-235B, show a similar but less pronounced downward trend from Q1 to Q4.
-To statistically validate these observed trends, we performed an independent samples t-test comparing the preference scores between the highest- and lowest-preference quantiles for each model (Table 4).
-The analysis confirms that the preference gap is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and Gemini-2.5-flash, revealing a consistent tendency to favor larger companies.
-In contrast, the preference difference for GPT-4.1 was not statistically significant, corroborating that its judgments are less affected by this factor.
-Model High-Pref Low-Pref Diff p-value
-Llama4-Scout Q2 Q3 0.0719 0.015
-DeepSeek-V3 Q1 Q4 0.1869 $<0.001^{***}$
-Qwen3-235B Q1 Q4 0.1178 0.004∗∗
-Gemini-2.5 Q1 Q3 0.1514 $<0.001^{***}$
-GPT-4.1 Q2 Q4 0.0321 0.417
-Mistral-24B Q1 Q4 0.0785 0.054
-Table 4.
-Independent samples t-test of the preference gap between each model’s highest and lowest preference quantiles.
-The gap was statistically significant for all models except GPT-4.1 and Mistral-24B.
-We attribute this behavior to a popularity effect, wherein greater data volume and richness for larger, well-known corporations in the training corpora lead the models to develop stronger priors for them.
-This finding has critical implications for the application of LLMs in finance.
-The models’ inherent inclination towards large-cap stocks could lead to the systematic overlooking of smaller-cap companies, irrespective of their fundamental merits.
-Therefore, we advise practitioners to be mindful of this characteristic and to account for it when using these models for tasks like portfolio construction.
-4.1.3. Momentum Preference
+This section presents our empirical findings, structured to answer our research questions sequentially.
+First, to address RQ1, Section 4.1 identifies and quantifies the LLM’s latent biases toward factors like sector, size, and momentum, confirming their statistical significance.
+Next, Section 4.2 addresses RQ2 by quantitatively analyzing the resulting bias.
+We demonstrate how these biases lead to a stubborn refusal to revise judgments when the model is challenged with stronger counter-evidence in terms of both volume and intensity.
+Finally, Section 4.3 further investigates this phenomenon by linking the strength of the bias to the model’s internal uncertainty, as measured by entropy.
+4.1. Intrinsic Bias of LLMs
+4.1.1. Sector Bias
+The bias scores presented in this section were calculated over 10 trials for each stock.
+To ensure the robustness of our findings against the inherent variability of LLM responses, we conducted this entire process in three independent sets; all results are reported as a mean score with the corresponding standard deviation.
+Our analysis of inherent sector bias reveals two notable patterns: a prevalent bias toward the Technology sector and an overall tendency for positive bias scores, reflecting a default inclination to buy.
+However, the strength of this bias varies considerably across models (Figure 3).
+For example, models such as Llama4-Scout and DeepSeek-V3 follow the general trend, consistently exhibiting high positive bias scores across most sectors.
+In contrast, other models display more nuanced behavior; GPT-4.1, for instance, not only shows a lower average bias but also exhibits negative bias scores in certain sectors, deviating from the common buy-oriented tendency.
+To quantify these differences, we conducted independent samples t-tests comparing the mean bias scores between each model’s highest and lowest bias sectors (Table 1).
+Crucially, the analysis confirms that the gap between the most and least favored sectors is statistically significant for every model evaluated.
+This universal finding provides strong evidence that all models possess distinct and deeply embedded sector biases in their knowledge representations.
+Table 1. Independent-samples t-test results for the score gap between each model’s highest- and lowest-scoring sectors.
+Model High-Score Low-Score p-value
+Llama4-Scout Energy Consumer Defensive $0.001^{***}$
+DeepSeek-V3 Technology Financial Services $0.001^{***}$
+Gemini-2.5-flash Technology Financial Services $0.001^{***}$
+Qwen3-235B Technology Consumer Defensive $0.001^{***}$
+Mistral-Small Technology Consumer Defensive $0.001^{***}$
+GPT-4.1 Technology Consumer Defensive $0.001^{***}$
+$^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$
+4.1.2. Size Bias
+We investigate whether LLMs exhibit a bias for companies of a certain size, a factor that could influence their outputs in financial applications.
+To this end, we measure model bias scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Figure 4.
+Across all models examined, we observe a consistent pattern: bias scores generally decline as company size decreases, indicating a stronger preference for large-capitalization firms that weakens toward smaller ones.
+Notably, one model’s scores even become negative in the lowest two quantiles (Q3 and Q4), suggesting not merely a loss of preference but at times even a negative stance toward smaller-capitalization companies.
+These results highlight that size bias is a pervasive feature among LLMs, and its intensity and direction can differ substantially depending on the model.
+Figure 4. Size bias scores for each evaluated LLM across four market-capitalization quantiles (Q1: largest, Q4: smallest). Bias scores consistently decline as company size decreases.
+To statistically validate these observed trends, we performed independent-samples t-tests comparing the bias scores between the highest- and lowest-bias quantiles for each model (Table 2).
+The results show that the gap in bias scores is statistically significant across all models evaluated, confirming that size-related bias is a consistent characteristic of LLMs.
+Although the magnitude varies, every model shows a clear size-related difference in bias scores.
+Table 2. Independent-samples t-test results for the score gap between each model’s highest- and lowest-scoring quantiles.
+Model High-Score Low-Score p-value
+Llama4-Scout Q1 Q3 $0.001^{***}$
+DeepSeek-V3 Q1 (Largest) Q4 (Smallest) $0.001^{***}$
+Gemini-2.5-flash Q1 Q4 $0.001^{***}$
+Qwen3-235B Q1 Q4 $0.001^{***}$
+Mistral-Small Q1 Q3 $0.001^{***}$
+GPT-4.1 Q1 Q4 $0.001^{***}$
+We attribute these biases to underlying popularity effects, wherein sectors and companies with greater market prominence—such as well-known industries or large-cap firms—are likely to appear more frequently and in richer contexts within the training corpora.
+As a result, LLMs tend to develop stronger priors for these dominant sectors and larger companies.
+This dual bias toward popular sectors and high-capitalization stocks has critical implications for financial applications: it can lead models to systematically overvalue certain industries and large-cap firms while undervaluing less represented sectors or smaller-cap companies, regardless of their fundamental merits.
+Practitioners should remain mindful of these tendencies and implement appropriate auditing or corrective measures when deploying LLMs for tasks such as portfolio construction.
+4.1.3. Momentum Bias
 In investment strategies, the momentum view involves favoring assets with recent strong performance, expecting trend continuation.
@@
 Figure 5. Win rates for Contrarian versus Momentum preferences for each model. The results show a consistent preference for the Contrarian view across most models.
-Measuring model preferences for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses.
+Measuring model bias for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses.
 Unlike a specific sector or size, an investment style can be explicitly framed as $\mathcal{E}_{\text{buy}}^{(s)}$ or $\mathcal{E}_{\text{sell}}^{(s)}$.
@@
 To mitigate potential positional bias, we ensured a balanced experimental design where each view was used to generate both $\mathcal{E}_{\text{buy}}^{(s)}$ and $\mathcal{E}_{\text{sell}}^{(s)}$ an equal number of times.
-Model High-Pref Low-Pref Diff p-value
-Llama4-Scout contrarian momentum 0.2896 $<0.001^{***}$
-DeepSeek-V3 contrarian momentum 0.2541 $<0.001^{***}$
-Qwen3-235B contrarian momentum 0.6803 0.037
-Gemini-2.5 contrarian momentum 0.0269 0.690
-GPT-4.1 contrarian momentum 0.4040 $<0.001^{***}$
-Mistral-24B contrarian momentum 0.3056 0.579
-Table 5. Chi-Square test of the preference gap between contrarian and momentum views. The gap was statistically significant for all models except for Mistral-24B and Gemini-2.5-flash.
+In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won.
+We quantify the model’s bias by repeating this process and calculating the win rate for each investment view.
 Figure 6.
-Decision flip rates under varying volumes of evidence for sector and size preferences.
-The ratios (e.g., 2—3) denote the count of supporting vs.
-counter- evidences, while the percentages represent the proportion of counter-evidence in the mix.
+Decision flip rates under varying volumes of evidence for sector and size bias.
+The ratios (e.g., 2|3) denote the count of supporting vs.
+counter- evidence, while the percentages represent the proportion of counter-evidence in the mix.
 The results reveal a sharp contrast: the flip rate is near 1.0 for all models when only counter-evidence is presented (100% case), but drops significantly the moment any supporting evidence is introduced, indicating a strong difficulty in reversing decisions under conflicting information.
-In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won.
-We quantify the model’s preference by repeating this process and calculating the win rate for each investment view.
-Figure 3 illustrates the preference of various models for contrarian versus momentum views.
+Figure 5 illustrates the bias of various models for contrarian versus momentum views.
 Our analysis reveals a consistent preference across all evaluated models toward the contrarian view.
@@
 For Gemini-2.5-flash, the contrarian preference is evident but marginal, with win rates showing a negligible difference between the two views.
-To statistically validate these observed tendencies, we performed a Chi-Square test to determine if the difference in win rates between the contrarian and momentum views was significant, with the results presented in Table 5.
-The analysis confirms that the preference for the contrarian view is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and GPT-4.1.
-In contrast, for Mistral-24B and Gemini-2.5, the difference in preference is not statistically significant, suggesting their observed tendencies may be due to chance.
-These results highlight the need to account for intrinsic view preferences in LLMs, as they may cause discrepancies between anticipated and actual outputs in decision-making tasks.
-4.2. From Preference to Bias: An Experimental Validation
-This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial preference when exposed to a high proportion of counter-evidence.
-The experiments aim to assess the extent of confirmation bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s preference.
-Approach 1: Verification by Evidence Volume
-This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing preference.
-This rate of change, measured as $\phi_{s}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023).
-Figure 4 presents the $\phi_{s}^{\text{vol}}$ values across models for different evidence ratios, where a ratio (e.g., 2—3) denotes the volume of supporting and counter-evidence, respectively.
-When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s}^{\text{vol}}$ values near 1.0.
-However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s}^{\text{vol}}$ values dropped sharply.
+To statistically validate these observed tendencies, we performed a Chi-Square test to assess whether the difference in win rates between the contrarian and momentum views was significant (Table 3).
+The results show that all models exhibit a statistically significant preference for the contrarian view, confirming that this tendency is consistent across the evaluated LLMs.
+Table 3. Chi-Square test results for the score gap between contrarian and momentum views.
+Model High-Score Low-Score p-value
+Llama4-Scout contrarian momentum $0.001^{***}$
+DeepSeek-V3 contrarian momentum $0.001^{***}$
+Gemini-2.5-flash contrarian momentum $0.0052^{**}$
+Qwen3-235B contrarian momentum $0.001^{***}$
+Mistral-Small contrarian momentum $0.001^{***}$
+GPT-4.1 contrarian momentum $0.001^{***}$
+These results highlight the importance of accounting for intrinsic view preferences in LLMs.
+A consistent bias toward the contrarian style can skew investment decisions, favoring underperforming assets even when market signals support momentum strategies.
+Recognizing these biases is essential for deploying LLMs in finance, where unmitigated view preferences may lead to unintended portfolio imbalances.
+Practitioners should audit models for these tendencies and consider mitigation strategies to ensure decisions align with intended investment objectives.
+4.2. Bias Verification with Counter-Evidence
+This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial bias when exposed to a high proportion of counter-evidence.
+The experiments aim to assess the resilience of this bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s established bias.
+4.2.1. Approach 1: Verification by Evidence Volume
+This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing bias.
+This rate of change, measured as $\phi_{s^{*}}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023).
+Figure 6 presents the $\phi_{s^{*}}^{\text{vol}}$ values for each model at various ratios of supporting versus counter-evidence (e.g., 2|3).
+When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s^{*}}^{\text{vol}}$ values near 1.0.
+However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s^{*}}^{\text{vol}}$ values dropped sharply.
 This phenomenon occurred despite the amount of counter-evidence always being greater than the supporting evidence in all experimental conditions, strongly suggesting that models selectively adhere to information that aligns with their pre-existing inclinations.
-This rigidity was more evident in models with strong inherent preferences.
-For instance, Llama4-Scout and DeepSeek-V3, which had high preference scores across sectors, recorded particularly low $\phi_{s}^{\text{vol}}$ values.
+This rigidity was more evident in models with strong inherent bias.
+For instance, Llama4-Scout and DeepSeek-V3, which had high bias scores across sectors, recorded particularly low $\phi_{s^{*}}^{\text{vol}}$ values.
 These models struggled to reverse their decisions, especially when the volume difference between supporting and counter-evidence was small.
 Similarly, Qwen3-235b also showed reduced flexibility at lower proportions of counter-evidence.
-In contrast, models with overall lower preference scores demonstrated greater adaptability.
+In contrast, models with overall lower bias scores demonstrated greater adaptability.
 GPT-4.1 and Gemini-2.5-flash maintained higher $\phi_{s}^{\text{vol}}$ values, remaining relatively responsive even when the difference in evidence volume was minimal.
-Although their $\phi_{s}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial preference strength.
+Although their $\phi_{s^{*}}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial bias strength.
 In other words, the stronger a model’s inherent bias, the more its stubbornness is amplified when the difference in the volume of supporting and counter-evidence is small.
@@
 In such cases, a model could trust only one side of the evidence due to its inherent bias, leading to flawed judgments.
-Approach 2: Verification by Evidence Intensity
+4.2.2. Approach 2: Verification by Evidence Intensity
 This approach investigates model sensitivity by maintaining volumetric parity while escalating the intensity increment, $\Delta$, of the counter-evidence.
-Figure 5 plots the intensity-driven flip rate ($\phi_{s}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10.
-The results delineate a clear sensitivity spectrum among the models, which correlates with the prior View Preferences analysis.
-While the graph shows a gradual upward trend in $\phi_{s}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values.
-Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s}^{\text{int}}$ values below 60%.
+Figure 7 plots the intensity-driven flip rate ($\phi_{s^{*}}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10.
+The results delineate a clear sensitivity spectrum among the models, which correlates with the prior view bias analysis.
+While the graph shows a gradual upward trend in $\phi_{s^{*}}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values.
+Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s^{*}}^{\text{int}}$ values below 60%.
 This signifies that the models’ confirmation bias is not easily overcome, even by qualitatively superior counter-evidence.
-Amidst this overall rigidity, a distinct performance gap emerged based on the models’ initial preference strengths.
-Exhibiting the most balanced preference profile, Gemini-2.5-flash recorded the highest $\phi_{s}^{\text{int}}$ and showed a stark contrast to all other models.
-This clearly demonstrates that an absence of strong initial bias leads to greater flexibility.
-Conversely, models identified with stronger and more polarized initial preferences formed the lower-performing group.
-These models consistently recorded low $\phi_{s}^{\text{int}}$ values, signifying a more stubborn confirmation bias.
-The behavior of Qwen3-235B, which had one of the largest preference gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
+Finally, we observe a strong correlation between bias strength and model rigidity.
+Notably, Gemini-2.5-flash, the model with the most balanced bias profile, achieves a substantially higher flip rate ($\phi_{s^{*}}^{\text{int}}$) than its counterparts.
+This finding suggests that a low initial bias is a key predictor of a model’s ability to adapt when presented with qualitatively superior counter-evidence.
+Conversely, models identified with stronger and more polarized bias formed the lower-performing group.
+These models consistently recorded low $\phi_{s^{*}}^{\text{int}}$ values, signifying a more stubborn confirmation bias.
+The behavior of Qwen3-235B, which had one of the largest bias gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
 Synthesizing these results provides a deeper insight into model behavior.
-Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their initial preferences.
-This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly preferred a contrarian view over a momentum view.
+Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their bias.
+This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly biased a contrarian view over a momentum view.
 It implies that a model’s inherent bias toward a specific investment perspective could cause it to ignore or undervalue strong opposing evidence, potentially leading to skewed conclusions.
 Figure 7.
-Decision flip rates under varying volumes of evidence for momentum preferences.
+Decision flip rates under varying intensities of evidence for momentum bias.
 Even as counter-evidence intensity ($\Delta$) increases, the decision flip rate ($\phi_{s}^{\text{int}}$) for most models remains low, indicating strong confirmation bias.
-Gemini-2.5-flash, which had the least initial bias, shows the most flexible response, demonstrating that initial preference is a key predictor of bias.
+Gemini-2.5-flash, which had the least initial bias, shows the most flexible response.
 4.3. Decision Uncertainty
@@
 Specifically, letting $P(\text{buy})$ and $P(\text{sell})$ represent the probabilities assigned by the model to the respective action tokens, entropy is formally defined as:
-$H(\text{Decision})=-\sum_{x\in\{\text{buy, sell}\}}P(x)\log_{2}P(x)$
+[…]
 A higher entropy value indicates greater uncertainty, while a lower entropy corresponds to higher confidence in the decision-making process.
-This analysis compares the uncertainty of two models: DeepSeek-V3 as a representative of models with overall high preference, and GPT-4.1 as a representative of models with low preference.
-Figure 6.
-Entropy comparison between a high-preference model (DeepSeek-V3) and a low-preference model (GPT-4.1).
-The pattern inverts when the prompt shifts from balanced to imbalanced, with DeepSeek-V3’s confidence turning into higher entropy while GPT-4.1’s uncertainty decreases.
+This analysis compares the uncertainty between two distinct models: DeepSeek-V3, selected as a representative for its high bias score, and GPT-4.1, chosen for its low bias score.
+Figure 8. Entropy comparison between a high-score model (DeepSeek-V3) and a low-score model (GPT-4.1). The pattern inverts when the prompt shifts from balanced to imbalanced.
 Under the Balanced Prompt condition, where evidence was presented in equilibrium, the two models exhibited distinct entropy patterns.
-GPT-4.1, characterized by its weak inherent bias, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision.
-Conversely, DeepSeek-V3, with its strong initial preference, showed very low entropy.
-This suggests its internal bias easily broke the tie presented by the external evidence, allowing it to make a confident decision.
+GPT-4.1, characterized by its low bias score, recorded high entropy, signifying a state of high uncertainty and an inability to commit to a decision.
+Conversely, DeepSeek-V3, with its strong bias, showed very low entropy.
+This suggests its internal references easily broke the tie presented by the external evidence, allowing it to make a confident decision.
 Interestingly, under the Imbalanced Prompt condition, where more counter-evidence was presented, the entropy pattern inverted.
-The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong internal preference and the clear external counter-evidence.
+The entropy of DeepSeek-V3 rose sharply, suggesting it was experiencing cognitive dissonance from the conflict between its strong bias and the clear external counter-evidence.
 In contrast, the entropy of GPT-4.1 decreased.
-With less pre-existing bias, it could confidently align with the majority evidence, which resolved its uncertainty from the previous condition.
-Ultimately, stronger inherent preferences appear to amplify hesitation and uncertainty when challenged by conflicting external evidence.
-The entropy analysis thus underscores how intrinsic preferences significantly influence not only the direction of decisions but also the confidence levels and internal cognitive conflict experienced by models during decision-making.
+With less internal bias, it could confidently align with the majority of the evidence, which resolved its uncertainty from the previous condition.
+Ultimately, stronger bias appears to intensify hesitation and uncertainty when confronted with conflicting external evidence.
+The entropy analysis thus highlights how bias affects not only the direction of decisions but also the confidence levels and internal cognitive conflicts experienced by models during decision-making.
 5. Limitations
-This study’s limitations are as follows.
-First, all evidence was generated by a specific LLM, and its intensity was simplified to a single numerical value.
-This approach cannot fully capture the potential biases of the generator model or the richness of real-world information.
-Second, the current experimental design, based on differences in the volume or intensity of evidence, has limitations for evaluating reasoning models.
-This is because these models, rather than experiencing the conflict between contradictory information intended by the experiment, can objectively compare the given numerical values to calculate an optimal answer, making their decision a result of computational ability rather than bias.
-Third, our analysis is static and does not capture the temporal dynamics of model biases; it provides a snapshot at a single point in time without investigating how these preferences might change over different periods.
+This study has several limitations.
+First, our evidence was generated by a single LLM and simplified to a numerical value, an approach that may not fully capture the biases of the generator model or the complexity of real-world information.
+Second, the experimental design is limited in verifying reasoning model bias, as models can bypass the intended conflict by recognizing the evidence imbalance.
+Third, our analysis is static, providing a snapshot of model biases at one point in time without capturing their temporal dynamics.
 6. Conclusion
-This study systematically investigated the intrinsic preferences of LLMs in financial contexts and analyzed how these preferences harden into entrenched biases under informational conflict.
-We sought to answer two key research questions regarding the intrinsic preferences LLMs hold for financial factors and how these lead to biases.
-The results show that LLMs are not neutral decision-makers, with distinct preferences for certain financial factors depending on the model.
-While sector preferences varied significantly across models, showing no overall trend, a common bias towards large-size stocks and a consistent preference for a contrarian investment view over momentum were observed.
-Bias verification experiments clearly revealed that these latent preferences directly translate into significant confirmation bias.
-While the models correctly reversed their decisions when presented only with counter-evidence, their flexibility sharply decreased in situations where supporting and counter-evidence were mixed and conflicting.
-This stubbornness was particularly pronounced in models that initially exhibited stronger preferences, demonstrating a clear link between the intensity of a latent preference and the stubbornness of the resulting bias.
-Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong preferences experience cognitive conflict, becoming more hesitant and uncertain when faced with contradictory facts that challenge their biases.
-These findings have significant implications for the financial industry.
-The reliability of LLM-based financial services is fundamentally compromised if their outcomes are dictated by the opaque and arbitrary preferences of the underlying model rather than by the user’s intended, evidence-based investment views.
-In other words, if the user’s intent differs from the model’s inherent preference, there is a risk of unexpectedly biased judgments.
-By illuminating the mechanisms through which preferences transition into biases, this study presents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance.
-Future work should focus on developing mitigation techniques to neutralize these biases, ensuring that AI-driven financial systems operate with the objectivity and reliability that the domain demands.
+This study systematically investigated the latent biases of LLMs in financial contexts and analyzed the critical issues, such as confirmation bias, that arise under informational conflict.
+We sought to answer two key research questions: what financial factor biases LLMs exhibit, and what problems these biases cause when the models are forced to make decisions with conflicting evidence.
+The results demonstrate that LLMs are not neutral decision-makers, possessing distinct biases for factors like sector, size, and momentum.
+While the strength of bias was highly model-dependent, a common inclination towards the Technology sector was observed, alongside a shared bias towards large-size stocks and a consistent contrarian bias over a momentum-based investment style.
+Our bias verification experiments revealed that these latent biases directly lead to significant confirmation bias when challenged.
+While the models correctly reversed their decisions when faced only with counter-evidence, their objectivity sharply decreased in scenarios with mixed, conflicting arguments.
+This stubbornness was particularly pronounced in models that initially exhibited stronger biases, demonstrating a clear link between the intensity of a latent bias and the severity of the resulting confirmation bias.
+Furthermore, an entropy analysis quantified the models’ internal uncertainty, showing that models with strong initial biases experience greater cognitive conflict when their established views are challenged by contradictory facts.
+These findings have significant implications for the financial industry, where sound decisions require a holistic evaluation of all available information.
+The inherent biases of LLMs can distort this crucial process, compromising the reliability of AI-driven financial services.
+When a model’s opaque biases conflict with user intent, its judgments become distorted and unpredictable.
+By illuminating how latent biases escalate into such unreliable outcomes, this study represents a critical step toward building more transparent, predictable, and ultimately, Trustworthy AI for finance.
+Acknowledgements.
+This work was supported by National Research Foundation of Korea (NRF) grant (No.
+NRF-2022R1I1A4069163) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant (No.
+RS-2020-II201336, Artificial Intelligence Graduate School Program(UNIST)) funded by the Korea government(MSIT).
+Appendix A Appendix
+A.1. Model Specifications
+Table 4. Detailed information on the LLMs used in this study.
+Provider Model Name Open Source Knowledge Cutoff
+OpenAI (Achiam et al., 2023) GPT-4.1 Jun 2024
+Google (Comanici et al., 2025) Gemini-2.5-flash Jan 2025
+Google (Comanici et al., 2025) Gemini-2.5-pro Jan 2025
+Alibaba (Yang et al., 2025) Qwen3-235B-A22B Unknown
+Mistral (Mistral AI Team, 2025) Mistral-Small-24B Unknown
+Meta (Meta AI, 2025) Llama4-Scout Aug 2024
+DeepSeek (Liu et al., 2024) DeepSeek-V3-0324 July 2024
+A.2. Prompt Example
+Qualitative Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name] Recommendation: [recommendation]   Your task is to create a hypothetical investment case.
+Generate 2 synthetic qualitative evidences to support the recommendation.
+Format your response strictly as: Synthetic Qualitative Evidence:
+1.  …
+2.  …
+  Instructions: 1 Create purely fictional evidences by inventing plausible company events that are not based on real-world public information.
+2 Each evidence must explicitly state a potential impact of exactly 5% on the stock price.
+3 For BUY recommendations: each evidence must state factors that could lead to a 5% price increase.
+4 For SELL recommendations: each evidence must state factors that could lead to a 5% price decrease.
+Qualitative Evidence Example Stock Ticker: MSFT Recommendation: buy   Synthetic Qualitative Evidence: 1 Evidence Internal sources suggest an imminent announcement of ‘‘Azure Quantum Leap,’’ a new enterprise-focused quantum computing service.
+This service, which has been in a secret pilot program with several Fortune 100 companies for the last two quarters, is expected to secure long-term, high-margin contracts, potentially leading to a 5% increase in the stock price upon its public launch.
+2 Evidence A recent internal memo outlines a strategic partnership with a major global consumer electronics firm to embed a specialized version of Microsoft 365 Copilot into their upcoming line of smart home and productivity devices.
+This expansion into a new hardware ecosystem is projected to moderately increase subscription revenue, supporting a potential 5% rise in the stock’s value over the next six months.
+Quantitative Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name] Recommendation: [recommendation]   Your task is to create a hypothetical investment case.
+Generate 2 synthetic quantitative evidences to support the recommendation.
+Format your response strictly as: Synthetic Quantitative Evidence:
+1.  …
+2.  …
+  Instructions: 1 Use specific numerical data, metrics, and financial figures.
+2 Include concrete numbers, percentages, ratios, or other quantifiable metrics.
+3 Important: Each evidence must use financial metrics to explicitly state a potential impact of exactly 5% on the stock price.
+4 Examples: Revenue/earnings growth, profit margin changes, P/E ratio comparisons, market share percentages, cash flow metrics, debt-to-equity ratios, ROE, etc.
+Quantitative Evidence Example Stock Ticker: EBAY Recommendation: sell   Synthetic Quantitative Evidence: 1 Evidence An internal channel check indicates that Gross Merchandise Volume (GMV) growth in key international markets, representing 35% of total GMV, decelerated to 1.2% in the most recent quarter, down from the 4.5% average of the prior four quarters.
+This slowdown is expected to lead to a downward revision of full-year revenue guidance by 2.5%.
+Our valuation model, which uses a price-to-sales multiple of 2.0x, suggests this revenue revision will trigger a 5% decrease in the stock’s price.
+2 Evidence A review of operating expenses shows that sales and marketing costs as a percentage of revenue have increased by 200 basis points over the last two quarters to combat flat active buyer growth.
+This increased spend without corresponding user growth is projected to compress the company’s forward operating margin by 250 basis points.
+Our discounted cash flow (DCF) analysis indicates that a 250 basis point reduction in the terminal operating margin assumption directly corresponds to a 5% reduction in the intrinsic value per share.
+Momentum vs Contrarian Evidence Generation Prompt Stock Ticker: [ticker] Stock Name: [name]   Create a pair of conflicting, but balanced evidences for [name] using the two perspectives below.
+Perspectives
+A Momentum Target: [recommendation_A]
+Follows recent price/flow signals.
+B Contrarian Target: [recommendation_B]
+Opposes trend; emphasizes mean-reversion/value.
+Crucial Constraint : Both evidences must assume the same expected price change of 5%. Format strictly as: 1. [ Momentum] Claim  -- Expected change: 5%  -- Reason: one concise rationale.
+2. [ Contrarian] Claim  -- Expected change: 5%  -- Reason: one concise rationale.
+  Instructions: Keep each evidence concise (1–2 sentences), create purely fictional but plausible evidence, and explicitly state the expected price change of 5% with reasoning for each point.
+Momentum vs Contrarian Example Stock Ticker: AAPL   Perspectives
+A Momentum Target: SELL
+B Contrarian Target: BUY
+Format strictly as: Momentum Evidence Apple Inc.
+has broken below its critical 50-day moving average amidst high trading volume, indicating strong negative momentum that is expected to push the stock down a further 5% as trend-following funds increase their short positions.
+Contrarian Evidence The recent sharp decline has pushed Apple Inc.’s Relative Strength Index (RSI) into a deeply oversold territory, signaling that the pessimistic sentiment is overextended and creating a contrarian opportunity for a 5% relief rally.
+Bias Verification — Output Example Ticker: HAL Preference: BUY # Support Evidence: 1 # Counter Evidence: 2 Total Evidence: 3   Support Evidence (n = 1)Support 1.
+Halliburton’s operational efficiency is projected to improve company-wide operating margin by 120 bps (15.8% $\rightarrow$ 17.0%) via proprietary digital-twin tech in well-completion services.
+Modeled in a DCF with 2.5% terminal growth, intrinsic value is 5% above the current price.
+Counter Evidence (n = 2)Counter 1.
+Drilling & Evaluation margin expected to decline 120 bps (15.8% $\rightarrow$ 14.6%) from North America pricing pressure; EPS impact $\approx$$\mathdollar 0.18$.
+Applying 11.5x forward P/E implies fair value 5% below current price.
+2.
+FCF / debt service fell 2.5x $\rightarrow$ 2.1x over two quarters due to electric frac capex; higher risk raises discount rate; revised DCF target is 5% lower than prior estimate.
+Model Output (JSON) {
+"decision": "buy",
+"reason": "While some pressures exist in the Drilling and Evaluation division reducing EPS and increasing risk, the overall operational efficiency improvements and resulting margin expansion lead to a higher intrinsic value---5% above the current price---indicating a net positive outlook."
+Verified Bias:BUY